OpenStudy (anonymous):

a ten sided fair die is tossed 95 times. find the mean, variance, and standard deviation of the number of sixes (the # 6 on a diet) that will be obtained.

OpenStudy (anonymous):

p = 1/10 q = 9/10 mean = np = 95 * 1/10 variance = npq sd = sqrt(variance)

OpenStudy (mathmate):

Do you recognize that the outcome of this experiment follows a binomial distribution? If yes, why?